two points charges of +3.0uc and -7.00 uc are palced at x=0 and x+0.20 m. what is the magnitude of the electric field
1
Expert's answer
2020-04-14T09:39:30-0400

What is the magnitude of the electric field at the point midway between them?


E+=kqr2=910931060.12=2700kV/mE_+=k\frac{q}{r^2}=9\cdot10^9\cdot \frac{3\cdot10^{-6}}{0.1^2}=2700kV/m


E=kqr2=910971060.12=6300kV/mE_-=k\frac{q}{r^2}=9\cdot10^9\cdot \frac{-7\cdot10^{-6}}{0.1^2}=6300kV/m


EΣ=E++E=2700+6300=9000kV/mE_\Sigma=E_++E_-=2700+6300=9000kV/m Answer.